# Panning for Gold: Expanding Domain-Specific Knowledge Graphs with General Knowledge

### v1

@@ 5.5. Relevant Entity Finding (RQ5)
In the DKGF process, accurate identification of general knowledge relevant to the DKG is a critical prerequisite to achieve reliable fusion. Failure to discover the correct cross-domain related entities results in erroneous fact generation and downstream inference. To address RQ5, we formulate and systematically study an essential yet under-explored task in the DKGF pipeline, namely Relevant Entity Finding (REF), which aims to identify a set of entities in the GKG that are semantically related to the DKG. We evaluated the performance of models on the REF task under varying training data conditions across two benchmark datasets. As shown in Table 5, ExeFuse consistently outperforms all representative benchmark configurations, especially in low-resource settings. This advantage stems from the executability-based relevance modeling in ExeFuse: instead of relying on surface-level semantic similarity, entity relevance is assessed through semantic operator mapping and latent program execution signals on the DKG, enabling the model to capture domain-relevant entities even when explicit semantic alignment is weak.
Furthermore, it is evident that the performance on the REF task exhibits a positive correlation with the performance on the DKGF task, as shown in Table 2. This indicates that REF is a key component in effectively solving the DKGF task.
@@ 5.6. Case study (RQ6)
To address RQ6 and provide an intuitive understanding of ExeFuse, we conduct a case study from both macro and micro perspectives. From a macro perspective, we analyze the prediction overlap of several benchmark configurations and visualize the results with the Venn diagram in Figure 4.
From the diagram we can find that ExeFuse has a significant portion of the proper predictions that do not intersect with several other benchmark configurations, which means that ExeFuse makes the right prediction on some test data that many other benchmark configurations predict incorrectly. This indicates that the fuzzy retriever and progressive KG fusion in ExeFuse play a crucial role in making accurate predictions. For a micro example, a newly fused fact from the test set, (Turki bin Faisal Al Saud, Host a visit, Brookings Institution, [2006-06]), was incorrectly predicted by DistMult-F. For the retrieved neighborhood fact (Paul A. Baran, employer, Brookings Institution, [-]), both Self-RAG-F and ChatEA-F still make erroneous predictions. Although the neighborhood information represents a correct fact, it does not directly contribute to the judgment of the current fusion fact involving Turki bin Faisal Al Saud and Brookings Institution. ExeFuse interprets general facts as executable semantic programs and combines execution evidence with structural consistency to ensure accurate fusion. As a result, ExeFuse outperforms other benchmark configurations on this new task.
Figure 4. The venn diagram of correct predictions across various DKGF benchmark configurations on DKGF(W-I)-S1. Intersections show shared same predictions.
@@ 6. Conclusion and Future Work
To the best of our knowledge, we are among the first to introduce a challenging and important new task, DKGF, to effectively enhance the coverage and structural integrity of DKGs by leveraging high-quality GKGs. To tackle the new task, we formally identified two key challenges (high ambiguity of domain relevance and cross-domain knowledge granularity misalignment) and developed a Fact-as-Program paradigm ExeFuse, which reformulates DKGF as executable semantic reasoning over domain-specific knowledge graphs. By mapping general facts to granularity-aware operators and verifying domain relevance through program executability, ExeFuse effectively integrates relevant and consistent knowledge from GKGs into DKGs. To address the scarcity of standardized benchmarks for DKGF, we construct two new benchmark datasets, DKGF(W-I) and DKGF(Y-I), and evaluate 21 representative configurations to provide the first standardized evaluation suite for DKGF. Extensive experiments demonstrate the effectiveness of ExeFuse and highlight the practical value of the new task.
DKGF represents an important yet underexplored research area with numerous valuable avenues for future investigation. In future work, we plan to explore several key directions: ① Developing more robust and reliable methods for domain relevance assessment to improve fusion accuracy; ② Investigating the critical role of temporal information in knowledge graph fusion; ③ Establishing comprehensive evaluation metrics specifically tailored for DKGF tasks that can better capture the quality and effectiveness of fused knowledge graphs.

### v2

@@ III. Design Choice Verification (Are choices optimal?)
5.3. Ablation Study (RQ3)
To rigorously evaluate the necessity of our architecture, we conducted a comprehensive ablation study on DKGF(W-I)-S1 (see Table 4). Specifically, we verify how each component contributes to addressing the two core DKGF challenges.
5.3.1. Justification of Complexity vs. Performance
We first address the concern of whether the neuro-symbolic complexity is warranted.
  • Single Affine Projection: We replaced the multi-rule logic execution with a single learnable linear transformation matrix. The performance drops significantly ($\Delta$-11.1%) to an F1 of 0.612. Notably, this result is comparable to the translation-based baseline TransE-F (F1 0.630) in Table 2. This confirms that simple linear mappings hit a performance bottleneck and are insufficient to bridge the complex semantic gap between GKG and DKG.
  • Direct Classification: Removing the reasoning module entirely leads to the worst performance ($\Delta$-15.3%), proving that the encoder alone cannot resolve the ambiguity of relevance without explicit logical transformation.
5.3.2. Impact of Key Modules & Challenge Resolution
We validate the contributions of each stage in the Fact-as-Program paradigm and their roles in solving specific tasks:
  • Latent Program Induction: Replacing the Structure-Aware Predicate ($\phi_{enc}$) with a standard linear encoder leads to the largest drop ($\Delta$-15.9%). This confirms that treating facts as structured programs, preserving the compositional logic of $(h,r,t)$, is the foundational basis for executable reasoning.
  • Neuro-Symbolic Execution (Addressing Challenge I): Removing the dynamic rule selector (using uniform weights) degrades performance by 9.2%, while removing execution entirely causes a 15.3% drop. This demonstrates that static similarity matching fails to capture implicit logical connections. The proposed neuro-symbolic execution effectively resolves Challenge I (Ambiguity of Domain Relevance) by inferring logical reachability beyond surface text.
  • Target Space Grounding (Addressing Challenge II): Removing the granularity calibration mechanism ($S_{exec}$) leads to a 12.7% drop. This finding is critical: it indicates that even if a fact is logically relevant, it may still fail to fuse correctly due to Challenge II (Granularity Misalignment). The grounding mechanism acts as a necessary compiler check, ensuring the transformed knowledge aligns with the specific density and granularity of the target domain.
5.3.3. Verification of Design Choices
Finally, we justify specific neuro-symbolic reasoning technical decisions:
  • Affine vs. MLP Rules: Replacing affine operators with non-linear MLPs degrades performance by 2.3%. We attribute this to the higher parameter count of MLPs leading to overfitting on the sparse DKG data, whereas affine transformations offer a robust and parameter-efficient representation for logic rules.
  • KDE vs. Min-Distance: Replacing KDE-based density estimation with simple Euclidean distance to the nearest centroid causes a 4.8% drop. This confirms that density-based verification is more reliable for quantifying granularity alignment than point-to-point distance metrics.
5.4. Efficiency Analysis (RQ4)
We evaluate ExeFuse’s efficiency against two categories: ① lightweight general-purpose and cross-task adaptation configurations (e.g., TransH-F, SimpleHHEA-F), which are efficient but limited in capturing complex domain interactions, and ② LLM-based general-purpose and cross-task adaptation configurations (e.g., ChatEA-F, Self-Consistency-F), which exploit LLMs but incur high computational overhead. As shown in Table 5 , ExeFuse achieves the best trade-off between effectiveness and efficiency: it delivers the highest accuracy (Avg. 0.685) while eliminating token consumption through our small-model paradigm, requiring only 42.6 seconds with substantial performance gains (17.9%-38.4%). This validates the advantage of our Fact-as-Program paradigm, which provides a new perspective for DKGF by reformulating knowledge fusion as executable neuro-symbolic reasoning without relying on expensive LLM API calls. These results validate its robustness in overcoming performance bottlenecks on the DKGF benchmark. The further analysis of theoretical complexity can be found in Section 4.6.
@@ 5.5. Relevant Entity Finding (RQ5)
In the DKGF process, accurate identification of general knowledge relevant to the DKG is a critical prerequisite to achieve reliable fusion. Failure to discover the correct cross-domain related entities results in erroneous fact generation and downstream inference. To address RQ5, we formulate and systematically study an essential yet under-explored task in the DKGF pipeline, namely Relevant Entity Finding (REF), which aims to identify a set of entities in the GKG that are semantically related to the DKG. We compare ExeFuse against embedding-based (DistMult-F) and LLM-based (Self-RAG-F, ChatEA-F) benchmark configurations. As shown in Table 6, we observe: (1) Resolving Ambiguity via Logical Reachability: ExeFuse significantly outperforms ChatEA-F (0.871 vs. 0.806 F1 on DKGF(W-I)-S1). While embeddings rely on surface similarity and LLMs lack structural constraints, ExeFuse redefines relevance as Logical Reachability. An entity is deemed relevant only if its associated fact successfully transforms into a valid DKG target state via neuro-symbolic execution. (2) Correlation with Fusion: Performance trends in REF closely mirror the main DKGF results in Table 6. This correlation validates our hypothesis that resolving domain relevance ambiguity (Stage 2) is the fundamental prerequisite for effective knowledge fusion.
Figure 3. The venn diagram of correct predictions on DKGF(W-I)-S1. Intersections show shared same predictions.
@@ 5.6. Case study (RQ6)
To intuitively demonstrate ExeFuse’s effectiveness, we conduct a multi-level analysis combining macro-statistics and micro-cases. (1) Macro Analysis: The Venn diagram in Figure 3 shows that ExeFuse correctly identifies a unique subset of fused facts ($27.9\%$) missed by both embedding-based and LLM-based baselines. This superiority stems from our neuro-symbolic approach, which captures complex logical dependencies and granularity alignments inaccessible to methods relying solely on surface text or static embeddings. (2) Micro Analysis: Consider the challenging inference: (Turki bin Faisal Al Saud, Host a visit, Brookings Institution). While DistMult-F fails due to low surface similarity, and LLMs (e.g., ChatEA-F) fail to bridge the semantic gap between the retrieved GKG fact (employer) and the target DKG relation (Host a visit), ExeFuse succeeds by treating the GKG fact as an executable program. Our Neuro-Symbolic Execution identifies the latent logic bridging these relations, while Target Space Grounding ensures the result aligns with the fine-grained granularity of political events. This confirms ExeFuse’s ability to resolve ambiguity through logical reachability.
@@ 6. Conclusion and Future Work
This paper introduces domain-Specific knowledge graph fusion (DKGF), a novel task aimed at enriching specialized knowledge graphs with general knowledge while overcoming the challenges of relevance ambiguity and granularity misalignment. To address these issues, we propose ExeFuse, a neuro-symbolic framework based on the Fact-as-Program paradigm. ExeFuse treats knowledge fusion as an executable logical process, employing neuro-symbolic execution to infer deep relevance and a target space grounding mechanism to calibrate knowledge granularity. We also establish the first standardized evaluation suite with two new benchmark datasets. Extensive experiments demonstrate that ExeFuse effectively bridges domain barriers to achieve superior fusion performance. In the future, we plan to enhance the framework by integrating LLMs to further improve the interpretability and reasoning depth of the neuro-symbolic execution process.